OpenStudy (shadyykatie):

A triangle has side lengths of 6, 8 and 10. Is it a right triangle? Explain.

OpenStudy (anonymous):

are you know the rule of the right triangle?

OpenStudy (shadyykatie):

Tbh i don't! @esamalaa

OpenStudy (anonymous):

\[a ^{2}+b ^{2}=c ^{2}\]

OpenStudy (shadyykatie):

@esamalaa i solved for A^2 + B^2 = C^2 but how would i know if it is a right triangle?

OpenStudy (anonymous):

let a=6 b=8 and then use the law if you get c^2=a^2+b^2 then this triangle is right

OpenStudy (whpalmer4):

The Pythagorean theorem applies only if it is a right triangle. If it is a right triangle, then \[a^2+b^2 = c^2\]If \(a^2+b^2 > c^2\), you have an acute triangle. If \(a^2+b^2 < c^2\), you have an obtuse triangle (or no triangle at all if \(a+b < c\))